Number of gm of oxygen in 32.2 g Na2SO4.10H2O is

Grade:

3 Answers

pankaj kumar raman
11 years ago

wt. of Na2SO4.10H2O=46+32+14*16+20=322gm

the above compound contain 14 gm Oxygen atom

hence 32.2 gm compound will contain 1.4 gm Oxygen atom

eaishwary bajpai
11 years ago

molecular mass of Na2SO4.10H2O =322 gm

322gm has 224 gm Oxgen 

 32.2 gm will have 22.4 gm oxygen


Rupsa Ghosh
6 years ago
No. of oxygen atoms in one molecule =4+10=14Thus moles of oxygen in n moles of compound=n*14Molar mass of Na2SO4.10H2O=46+32+64+180=322g/molThus 32.2g=32.2/322=0.1 mols Thus moles of O=0.1*14=1.4So weight is 16*1.4=22.4g